You have a battery marked "6.00 V." When you draw a current of 0.347 A from it, the potential difference between its terminals is 5.39 V. What is the potential difference when you draw 0.583 A?
Answer #1

V = E - I r

5.39 = 6 - (0.347 r)

r = 1.76 ohm


when I = 0.583 A


V = 6 - (0.583 x 1.76)

V =4.98 Volt ......Ans